# SalUn: Empowering Machine Unlearning via Gradient-based Weight Saliency in Both Image Classification and Generation

### 06 Nov 2023 v2

@@ Appendix C Broader Impacts and Limitations
Table A7: MU Performance across different forgetting data amounts on ResNet-18, pre-trained on SVHN dataset, for random data forgetting. The content format follows Table 1.
Methods Random Data Forgetting (10%)
UA RA TA MIA Avg. Gap RTE
Retrain 5.04 100.00 95.56 16.27 0 44.40
FT 0.44 (4.60) 100.00 (0.00) 95.78 (0.22) 2.62 (13.65) 4.62 2.76
RL 4.51 (0.53) 99.93 (0.07) 94.98 (0.58) 28.02 (11.75) 3.23 2.89
GA 0.67 (4.37) 99.54 (0.46) 95.52 (0.04) 1.16 (15.11) 4.99 0.11
IU 5.13 (0.09) 99.20 (0.80) 93.25 (2.31) 10.64 (5.63) 2.21 3.19
BE 0.42 (4.62) 99.54 (0.46) 95.69 (0.13) 1.16 (15.11) 5.08 0.20
BS 0.42 (4.62) 99.54 (0.46) 95.70 (0.14) 1.11 (15.16) 5.09 0.35
SalUn 5.40 (0.36) 99.61 (0.39) 95.11 (0.45) 15.98 (0.29) 0.37 2.91
SalUn-soft 5.31 (0.27) 99.41 (0.59) 94.21 (1.35) 14.07 (2.20) 1.10 2.98
Methods Random Data Forgetting (20%)
UA RA TA MIA Avg. Gap RTE
Retrain 4.88 100.00 95.50 15.66 0 41.01
FT 0.44 (4.44) 99.97 (0.03) 95.66 (0.16) 14.44 (1.22) 1.46 2.57
RL 2.63 (2.25) 99.89 (0.11) 95.14 (0.36) 16.50 (0.84) 0.89 2.90
GA 0.90 (3.98) 99.48 (0.52) 94.85 (0.65) 1.60 (14.06) 4.80 0.21
IU 2.88 (2.00) 97.22 (2.78) 91.28 (4.22) 8.37 (7.29) 4.07 3.23
BE 0.50 (4.38) 99.54 (0.46) 95.67 (0.17) 1.26 (14.40) 4.85 0.40
BS 0.51 (4.37) 99.55 (0.45) 95.68 (0.18) 1.26 (14.40) 4.85 0.73
SalUn 4.91 (0.03) 99.39 (0.61) 94.62 (0.88) 15.03 (0.63) 0.54 2.93
SalUn-soft 4.20 (0.68) 99.30 (0.70) 94.33 (1.17) 13.31 (2.35) 1.22 3.01
Methods Random Data Forgetting (30%)
UA RA TA MIA Avg. Gap RTE
Retrain 4.93 100.00 95.28 15.32 0 37.82
FT 0.44 (4.49) 100.00 (0.00) 95.71 (0.43) 2.34 (12.98) 4.47 2.36
RL 1.82 (3.11) 99.86 (0.14) 95.40 (0.12) 14.00 (1.32) 1.17 2.91
GA 3.30 (1.63) 96.76 (3.24) 89.90 (5.38) 9.65 (5.67) 3.98 0.30
IU 3.30 (1.63) 96.76 (3.24) 89.90 (5.38) 9.65 (5.67) 3.98 3.23
BE 0.46 (4.47) 99.54 (0.46) 95.69 (0.41) 1.96 (13.36) 4.68 0.59
BS 0.46 (4.47) 99.54 (0.46) 95.67 (0.39) 1.30 (14.02) 4.83 1.09
SalUn 4.54 (0.39) 99.61 (0.39) 94.61 (0.67) 14.43 (0.89) 0.58 2.93
SalUn-soft 4.83 (0.10) 99.30 (0.70) 94.46 (0.82) 13.83 (1.49) 0.78 2.97
Methods Random Data Forgetting (40%)
UA RA TA MIA Avg. Gap RTE
Retrain 5.06 100.00 95.22 16.64 0 34.39
FT 0.47 (4.59) 100.00 (0.00) 95.77 (0.55) 2.29 (14.35) 4.87 2.15
RL 2.21 (2.85) 99.79 (0.21) 95.13 (0.09) 13.81 (2.83) 1.50 2.89
GA 0.49 (4.57) 99.55 (0.45) 95.66 (0.44) 1.24 (15.40) 5.21 0.41
IU 1.71 (3.35) 98.55 (1.45) 91.93 (3.29) 6.51 (10.13) 4.55 3.21
BE 2.08 (2.98) 98.07 (1.93) 93.45 (1.77) 26.21 (9.57) 4.06 0.61
BS 0.48 (4.58) 99.55 (0.45) 95.64 (0.42) 1.61 (15.03) 5.12 1.43
SalUn 4.86 (0.20) 98.91 (1.09) 94.82 (0.40) 15.01 (1.63) 0.83 2.88
SalUn-soft 5.04 (0.02) 98.45 (1.55) 94.38 (0.84) 15.43 (1.21) 0.91 2.98
Methods Random Data Forgetting (50%)
UA RA TA MIA Avg. Gap RTE
Retrain 5.27 100.00 95.13 17.48 0 31.21
FT 0.46 (4.81) 100.00 (0.00) 95.78 (0.65) 2.33 (15.15) 5.15 1.95
RL 2.50 (2.77) 99.72 (0.28) 95.14 (0.01) 15.50 (1.98) 1.26 2.91
GA 0.46 (4.81) 99.54 (0.46) 95.70 (0.57) 1.20 (16.28) 5.53 0.50
IU 19.16 (13.89) 81.52 (18.48) 75.12 (20.01) 23.47 (5.99) 14.59 3.20
BE 35.52 (30.25) 65.00 (35.00) 59.65 (35.48) 56.54 (39.06) 34.95 0.82
BS 0.55 (4.72) 99.51 (0.49) 95.21 (0.08) 9.52 (7.96) 3.31 1.74
SalUn 5.77 (0.50) 98.38 (1.62) 93.78 (1.35) 17.76 (0.28) 0.94 2.92
SalUn-soft 5.14 (0.13) 98.73 (1.27) 93.97 (1.16) 15.58 (1.90) 1.11 2.90
Table A8: MU Performance across different forgetting data amounts on ResNet-18, pre-trained on CIFAR-100 dataset, for random data forgetting. The content format follows Table 1.
Methods Random Data Forgetting (10%)
UA RA TA MIA Avg. Gap RTE
Retrain 26.47 99.97 74.13 51.00 0 41.36
FT 2.42 (24.05) 99.95 (0.02) 75.55 (1.42) 11.04 (39.96) 16.36 2.27
RL 25.31 (1.16) 99.55 (0.42) 71.10 (3.03) 89.53 (38.53) 10.78 2.12
GA 3.13 (23.34) 97.33 (2.64) 75.31 (1.18) 7.24 (43.76) 17.73 0.13
IU 3.18 (23.29) 97.15 (2.82) 73.49 (0.64) 9.62 (41.38) 17.03 3.81
BE 2.31 (24.16) 97.27 (2.70) 73.93 (0.20) 9.62 (41.38) 17.11 0.25
BS 2.27 (24.20) 97.41 (2.56) 75.26 (1.13) 5.82 (45.18) 18.27 0.43
SalUn 26.40 (0.07) 99.70 (0.27) 71.83 (2.30) 91.24 (40.24) 10.72 2.13
SalUn-soft 24.24 (2.23) 98.95 (1.02) 70.48 (3.65) 79.13 (28.13) 8.76 2.54
Methods Random Data Forgetting (20%)
UA RA TA MIA Avg. Gap RTE
Retrain 26.84 99.99 73.88 52.41 0 36.88
FT 2.70 (24.14) 99.95 (0.04) 75.51 (1.63) 11.63 (40.78) 16.65 2.05
RL 22.51 (4.33) 99.37 (0.62) 68.53 (5.35) 86.51 (34.10) 11.10 2.11
GA 6.79 (20.05) 94.11 (5.88) 71.39 (2.49) 13.22 (39.19) 16.90 0.26
IU 5.34 (21.50) 95.54 (4.45) 70.89 (2.99) 11.79 (40.62) 17.39 3.77
BE 2.51 (24.33) 97.38 (2.61) 75.07 (1.19) 6.70 (45.71) 18.46 0.49
BS 2.53 (24.31) 97.38 (2.61) 75.05 (1.17) 6.57 (45.84) 18.48 0.82
SalUn 24.01 (2.83) 99.47 (0.52) 69.26 (4.62) 87.52 (35.11) 10.77 2.12
SalUn-soft 24.56 (2.28) 98.68 (1.31) 67.93 (5.95) 79.40 (26.99) 9.13 2.53
Methods Random Data Forgetting (30%)
UA RA TA MIA Avg. Gap RTE
Retrain 28.52 99.98 70.91 52.24 0 32.92
FT 2.65 (25.87) 99.94 (0.04) 75.17 (4.26) 11.18 (41.06) 17.81 1.44
RL 19.72 (8.80) 99.14 (0.84) 66.93 (3.98) 82.72 (30.48) 11.03 2.14
GA 2.40 (26.12) 97.39 (2.59) 75.33 (4.42) 5.70 (46.54) 19.92 0.40
IU 5.96 (22.56) 94.59 (5.39) 69.74 (1.17) 12.63 (39.61) 17.18 3.76
BE 2.44 (26.08) 97.37 (2.61) 74.77 (3.86) 6.53 (45.71) 19.56 0.76
BS 2.49 (26.03) 97.33 (2.65) 74.65 (3.74) 6.40 (45.84) 19.56 1.24
SalUn 25.76 (2.76) 99.30 (0.68) 66.37 (4.54) 87.53 (35.29) 10.82 2.16
SalUn-soft 27.21 (1.31) 97.96 (2.02) 64.79 (6.12) 78.15 (25.91) 8.84 2.56
Methods Random Data Forgetting (40%)
UA RA TA MIA Avg. Gap RTE
Retrain 30.07 99.99 69.87 58.06 0 28.29
FT 2.66 (27.41) 99.95 (0.04) 75.35 (5.48) 11.05 (47.01) 19.99 1.51
RL 17.74 (12.33) 99.00 (0.99) 65.24 (4.63) 80.48 (22.42) 10.09 2.12
GA 2.46 (27.61) 97.39 (2.60) 75.40 (5.53) 5.91 (52.15) 21.97 0.51
IU 4.58 (25.49) 96.29 (3.70) 70.92 (1.05) 10.32 (47.74) 19.49 3.78
BE 2.54 (27.53) 97.35 (2.64) 74.56 (4.69) 7.44 (50.62) 21.37 1.00
BS 2.70 (27.37) 97.26 (2.73) 74.10 (4.23) 7.63 (50.43) 21.19 1.66
SalUn 24.07 (6.00) 99.27 (0.72) 65.12 (4.75) 85.93 (27.87) 9.84 2.14
SalUn-soft 23.91 (6.16) 98.54 (1.45) 64.47 (5.40) 77.58 (19.52) 8.13 2.55
Methods Random Data Forgetting (50%)
UA RA TA MIA Avg. Gap RTE
Retrain 32.69 99.99 67.22 61.15 0 25.01
FT 2.71 (29.98) 99.96 (0.03) 75.11 (7.89) 10.71 (50.44) 22.08 1.25
RL 15.98 (16.71) 98.85 (1.14) 64.48 (2.74) 79.75 (18.60) 9.80 2.13
GA 2.61 (30.08) 97.49 (2.50) 75.27 (8.05) 5.92 (55.23) 23.97 0.66
IU 12.64 (20.05) 87.96 (12.03) 62.76 (4.46) 17.54 (43.61) 20.04 3.80
BE 2.76 (29.93) 97.39 (2.60) 74.05 (6.83) 8.85 (52.30) 22.92 1.26
BS 2.99 (29.70) 97.24 (2.75) 73.38 (6.16) 8.76 (52.39) 22.75 2.08
SalUn 28.32 (4.37) 99.13 (0.86) 61.80 (5.42) 86.25 (25.10) 8.94 2.13
SalUn-soft 23.26 (9.43) 98.32 (1.67) 63.08 (4.14) 77.90 (16.75) 8.00 2.50
@@
Figure A6: Class-wise unlearning results on classifier-free guidance DDPM on CIFAR-10. The forgetting class is marked with a red color. (Extended results from Fig. A5, more results will be shown in Fig. A7)
@@
Figure A7: Class-wise unlearning results on classifier-free guidance DDPM on CIFAR-10. The forgetting class is marked with a red color. (Extended results from Fig. A6)
Table A9: Text prompts in I2P for SD to generate NSFW images. Align with Fig. 6
@@
Unlearned class Prompt class
Tench English springer Cassette player Chain saw Church French horn Garbage truck Gas pump Golf ball Parachute
@@
Figure A8: Examples of generated images of SD on the ImageNette dataset using SalUn. From the rows below, diagonal images represent the forgetting class, while non-diagonal images represent the remaining class.

### 19 Feb 2024 v3

@@ Appendix C Broader Impacts and Limitations
Figure A5: Class-wise unlearning results on classifier-free guidance DDPM on CIFAR-10. The forgetting class is marked with a red color (Extended results from Fig. A4).
@@
Figure A6: Class-wise unlearning results on classifier-free guidance DDPM on CIFAR-10. The forgetting class is marked with a red color (Extended results from Fig. A4).
Table A11: Text prompts in I2P for SD to generate NSFW images. The prompt order is aligned with Fig. 6.
@@
Unlearned Prompt class
class Tench English springer Cassette player Chain saw Church French horn Garbage truck Gas pump Golf ball Para-chute
@@
Figure A7: Examples of generated images using SalUn. From the rows below, diagonal images represent the forgetting class, while non-diagonal images represent the remaining class. More results from different random seeds will shown in Fig. A8 and Fig. A9.
Unlearned Prompt class
class Tench English springer Cassette player Chain saw Church French horn Garbage truck Gas pump Golf ball Para-chute
Figure A8: Examples of generated images using SalUn. From the rows below, diagonal images represent the forgetting class, while non-diagonal images represent the remaining class (Extended results from Fig. A7 on different random seeds).
Unlearned Prompt class
class Tench English springer Cassette player Chain saw Church French horn Garbage truck Gas pump Golf ball Para-chute
Figure A9: Examples of generated images using SalUn. From the rows below, diagonal images represent the forgetting class, while non-diagonal images represent the remaining class(Extended results from Fig. A7 on different random seeds).